# Gravitational waves and black holes from the phase transition in models of dynamical symmetry breaking

### v1

@@ 1 Introduction
In this work we analyze the cosmological implications of such BSM phase transitions. Besides an enhanced GW signal, a supercooled FOPT also leads to the formation of primordial black holes (PBHs) via a mechanism suggested in [58, 59, 60, 61, 62, 63] that recently received interest [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. FOPTs proceed via the nucleation of bubbles of the broken phase in an initial background of the symmetric phase [86, 87, 88]. In a supercooled regime, the energy density of the Universe in the symmetric phase is dominated by the vacuum energy which acts as a cosmological constant and leads to an inflationary period. In a nucleated bubble, instead, such energy is quickly converted into radiation and the corresponding patch expands slower. Since bubble nucleation is a stochastic phenomenon, and since, in a supercooled regime, regions outside the nucleated bubbles dilute faster than those inside, a region where nucleation happens later becomes over-dense and eventually collapses, if large enough, into a PBH.
@@ 2.3 Gravitational waves from the phase transition
where $c_{l}=0.064$ and $c_{h}=0.48$. $S_{\rm col}$ equals unity at the peak frequency $f_{\text{col}}$
$f_{\text{col}}=1.65\times 10^{-5}\,{\rm Hz}~{}\,\Xi_{\text{col}}(v_{\text{wall}})\,\frac{\beta}{H}\,\frac{T_{\rm nuc}}{100\,{\rm GeV}}\,\left(\frac{g_{\star}}{100}\right)^{1/6}\ ,$ (2.22)
@@ 3.2 General parameter space of the model
Fig. 5 shows contour levels of the nucleation temperature and of the inverse duration $\beta/H$ of the phase transition. Interesting values are obtained in a region ranging from too large non-perturbative couplings down to too small couplings, such that the phase transition is ended by QCD rather than by nucleation. Fig. 5 shows, in the same plane, contour values of the PBH masses. Fig. 5 shows contour values of the PBH abundance. As it is exponentially sensitive to parameter values, the figures focuses on a small portion of the parameter space where the PBH abundance is around the DM abundance.
@@
Fig. 6 also shows that significant GW signals arise in the white allowed region (see also [162]). Each future GW experiment would probe a band in parameter space. More in detail, fig. 7 shows how the various possible GW experiments cover the parameter space, by being sensitive to different frequencies. As a specific example, fig. 8 shows that the expected Signal-to-Noise Ratio at LISA can reach large values. However, we here neglected the astrophysical foregrounds, expected with the spectrum illustrated in fig. 1. It’s difficult to anticipate how much foregrounds will limit the sensitivity to primordial GW, as some foregrounds can be partially subtracted.
@@ 3.3 Particles as DM
Figure 6: Summary plot for the minimal model. The right panel assumes a larger mixed quartic $\lambda_{HS}$, while in the left panel $\lambda_{HS}$ is so small that its running leads to multi-phase effects. The cosmological relic abundance of particle (of Primordial Black Hole) DM is over-abundant above the green (red) curve. The shaded regions are excluded by DM direct detection (purple); Higgs $h\to ss$ decays (beige); bounds on Higgs/dilaton mixing (orange); too large quartic couplings (pink).
@@ 4 Model with SU(2) vector DM
The cosmological DM abundance is again reproduced when $\sigma_{0}\approx 1/(23\,{\rm TeV})^{2}$. We consider the regime where $g_{X}\sim 1$ is large enough that super-cooling is ended by nucleation, rather than by QCD. Then $T_{\rm RH}=(136/64g_{*})^{1/4}M_{X}/\pi\approx M_{X}/8.5$ is again larger than the DM decoupling temperature.
@@
Figure 8: Expected Signal-Noise Ratio at LISA as function of the dilaton $m_{s}$ for different fixed values of the DM mass $m_{s^{\prime}}$, and neglecting astrophysical foregrounds. A SNR $\gtrsim 10$ allows GW detection.

### v2

@@ 1 Introduction
In this work we analyze the cosmological implications of such BSM phase transitions. Besides an enhanced GW signal, a supercooled FOPT also leads to the formation of primordial black holes (PBHs) via different mechanisms. Bubble collisions have been studied e.g. in [58]. We focus on a mechanism suggested in [59, 60, 61, 62, 63, 64] that recently received interest [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]. FOPTs proceed via the nucleation of bubbles of the broken phase in an initial background of the symmetric phase [88, 89, 90]. In a supercooled regime, the energy density of the Universe in the symmetric phase is dominated by the vacuum energy which acts as a cosmological constant and leads to an inflationary period. In a nucleated bubble, instead, such energy is quickly converted into radiation and the corresponding patch expands slower. Since bubble nucleation is a stochastic phenomenon, and since, in a supercooled regime, regions outside the nucleated bubbles dilute faster than those inside, a region where nucleation happens later becomes over-dense and eventually collapses, if large enough, into a PBH.
@@ 2.3 Gravitational waves from the phase transition
where $c_{l}=0.064$ and $c_{h}=0.48$. $S_{\rm col}$ equals unity at the peak frequency $f_{\text{col}}$ given by
$f_{\text{col}}=1.65\times 10^{-5}\,{\rm Hz}~{}\,\Xi_{\text{col}}(v_{\text{wall}})\,\frac{\beta}{H}\,\frac{T_{\rm RH}}{100\,{\rm GeV}}\,\left(\frac{g_{\star}}{100}\right)^{1/6}\qquad\hbox{for}\qquad T_{\rm RH}\gg T_{\rm nuc}$ (2.22)
@@ 3.2 General parameter space of the model
Fig. 5 shows contour levels of the nucleation temperature and of the inverse duration $\beta/H$ of the phase transition. Interesting values are obtained in a region ranging from too large non-perturbative couplings down to too small couplings, where the phase transition is ended in a too rapid way by QCD rather than by nucleation [81, 164]. Fig. 5 shows, in the same plane, contour values of the PBH masses. Fig. 5 shows contour values of the PBH abundance. As it is exponentially sensitive to parameter values, the figures focuses on a small portion of the parameter space where the PBH abundance is around the DM abundance.
@@
Fig. 6 also shows that significant GW signals arise in the white allowed region (see also [165]). Each future GW experiment would probe the region in parameter space above the indicated curve, up to the allowed boundary where $f_{\rm PBH}\approx 1$. We do not study the region of parameter space where QCD (rather than nucleation) ends the phase transition. More in detail, fig. 7 shows how the various possible GW experiments cover the parameter space, by being sensitive to different frequencies. As a specific example, fig. 8 shows that the expected Signal-to-Noise Ratio at LISA can reach large values. However, we here neglected the astrophysical foregrounds, expected with the spectrum illustrated in fig. 1. It’s difficult to anticipate how much foregrounds will limit the sensitivity to primordial GW, as some foregrounds can be partially subtracted.
@@ 3.3 Particles as DM
Figure 6: Summary plot for the minimal model. The right panel assumes a larger mixed quartic $\lambda_{HS}$, while in the left panel $\lambda_{HS}$ is so small that its running leads to multi-phase effects. The cosmological relic abundance of particle (of Primordial Black Hole) DM is over-abundant above the green (red) curve. Gravitational waves appear detectable above the curves indicated as LISA and ET. The shaded regions are excluded by DM direct detection (purple); Higgs $h\to ss$ decays (beige); bounds on Higgs/dilaton mixing (orange); too large quartic couplings (pink).
@@ 4 Model with SU(2) vector DM
The cosmological DM abundance is again reproduced when $\sigma_{0}\approx 1/(23\,{\rm TeV})^{2}$. We restrict our analysis to the regime where $g_{X}\sim 1$ is large enough that super-cooling is ended by nucleation, rather than by QCD. Then $T_{\rm RH}=(136/64g_{*})^{1/4}M_{X}/\pi\approx M_{X}/8.5$ is again larger than the DM decoupling temperature.
@@
The phase transition in the non-scale-invariant version of the model was studied in [167].
Figure 8: Expected Signal-Noise Ratio. at LISA as function of the dilaton $m_{s}$ for different fixed values of the DM mass $m_{s^{\prime}}$, and neglecting astrophysical foregrounds. A SNR $\gtrsim 10$ allows GW detection.